OpenStudy (anonymous):

A snake is slithering toward you at 1.5 m/s. If you start walking when it is 5.0 m away, how fast must you go in order that the snake not overtake you when you have gone 40.0 m?

OpenStudy (stormfire1):

First, since the snake is travelling at a constant velocity, find out how long it takes the snake to go 45m (you have a 5m head start). \[\frac{45m}{1.5m/s} = 30s\] Now you just need to go your 40 meters in that amount of time: \[Xf−Xi=1/2(Vi+Vf)t\] \[40m−0m=1/2(0+Vf)(30s)\]\[\frac{40m∗2}{30s} =Vf=2 \frac{2}{3} m/s\] Since you're starting from rest, your average speed must be half of that which is \[1\frac{1}{3}m/s\] So at 40m, you and the snake will be at the same point however you will be going \[2 \frac{2}{3}m/s\] at that point so the snake cannot overtake you. Looking back at the question, I believe it wants to know what your average speed needs to be...which is:\[>1\frac{1}{3}m/s\]